INAF queue observing, 2018 Feb 01-02

Observer: F. Cusano, A. Gargiulo
Telescope Operator: D. Gonzalez-Huerta
Support Astronomer: B. Rothberg
Instrument: LUCI1

Summary

At the sunset sky is mostly clear with some high cirrus
instruments: the 31st Jan the DX-AdSec was reinstalled and now the right side need to be calibrated, so we can not use LUCI2 and MODS2
LBC works, LUCI1 has some problem with the tip-tilt mirror, but AO should work

Night duration (12\xB0 el twilight) 2:00am- 13:15pm UT = 11 h 15m

Observing time: 9h 45m (87%)

Weather time loss: 0h (0%)

technical time loss 1 30h (13%)

Data Summary

Project_RUN Objects Instrument Filter ExpTime Notes
Nisini, imaging Field3 LUCI1 H2, FeII, J 0.3h completed
Prandoni USS188 LUCI1 K, J 0.6h completed
Vanzella JDz11 LUCI1 Ks 0.75h some clouds at the end
Prandoni USS36 LUCI1 K 0.4h  
Saracco C1_23152 LUCI1 200HK + HKspec 2h  
Prandoni USS172 LUCI1 K, J 0.6h completed
Prandoni USS202 LUCI1 K, J 0.6h completed
LUCI cal flat arc LUCI1 200g HKspec   Saracco

Details

01:35 slewing to imaging Nisini Field3, ima_field3_luci1.xml
  • 01:36 seeing 0.6" on guider
  • first image luci1.20180202.0001.fits
  • 02:03 seeing 0.7" on guider

2:13 slewing to Prandoni USS188 imaging K
  • no guide star found, checking poitnin
  • 2;24 we change the guide star, now is ok
  • 2:27 still trying to collimate, the TO sending manual corrections to the primary
  • 2:31 the primary converged, starting integration, first image luci1.20180202.0028.fits
  • 2:31 seeing 0.6" on guider

03:04 sending Prandoni USS188 imaging J
  • first image luci1.20180202.0048.fits
  • 3:05 seeing 0.5" on guider

3:25 slewing to Vanzella AO program
  • 3:27 problem in moving the N30 camera, 3:32 problem solved by the SA
  • 3:29 the N30 field stop mask is in, so the movement of the masks should be fine
  • 3:34 seeing 0.6" on guider
  • bin 2, freq 400
  • NCPA OFF
  • 3:40, we stop the integration, the N30 camera is out of alignment, the SA is checking and aligning, 62-65 is an alignment cal image
  • 3:49 start integration again, first good image 66
  • 3:52 seeing 0.7" on guider
  • average fwhm of 0.15-0.2" at 4:00
  • 4:18* completed the first cicle sending again the script at 4:19
  • 4:20 seeing 0.7" on guider
  • 4:42 average fwhm 0.1-0.2", seeing 0.7" on guider, 0.6" dimm
  • 4:52 completed the second cicle sending again the script at 4:53
  • 4:54 seeing 0.7" on guider
  • 4:55 loop opened the AO support is checking
  • 4:57 running again the script
  • 5:05 problem with the offset, pause and then re-sending the script
  • loop opens images 126 not good
  • 5:17 loop closed, sending again the script
  • 5:18 problem with the offset stopping the script, loop opened
  • 5:25 sending again the script
  • 5:28 loop opens paused the script, some clouds passing and reflection of the moon disturbs the ao observations
  • we finish the AO observations

5:38 slew to Prandoni USS36 imaging K
  • 5:41 seeing 1" on guider
  • 5:51 seeing 1.6" on guider, 1.1 on dim
  • seeing jumped to 2.2 at 5:56
  • 5:57 paused the script, the TO is checking the collimation
  • 6:07 collimation problem solved we resumed the script, seeing 1.1" on guider
  • 6:10 seeing 1" on guider
  • 6:18 jump in the seeing at 2.1"
  • checking the collimation, clear active optics
  • 7:02 problem solved, seeing 1" on guider now
  • 7:03 sending again the K script
  • 7:31 seeing 0.9" on guider but still some problem with the collimations (astigmatism)

7:35 slewing to telluric of Saracco, TelluricLuci1.xml, HIP 56147
  • 7:37 no guide star found, the TO is checking the pointing
  • 7:41 send again the script, still problem, check again the pointing
  • 7:45 send again the scirp
  • 7:46 change the guide star, we send the new script
  • acquisition 170-172
  • spectra 173-175
  • 7:56 seeing 0.7 on guider

07:59 slewing to science of Saracco, Target C1_23152
  • no guise star found, check the pointing
  • 08:07 send again the preset, now is ok
  • acq 176-180
  • 8:22 seeing 0.6" on guider
  • spec 181-192
  • 9:14 seeing 0.6"

09:33 slewing to Gilli J science
  • seeing 0.7" on guider
  • target too close to the moon, the guide star is very faint and the telescope can not gude, we move back to Saracco

09:46 slew to Saracco LBT-2017B-C2743_3.xml,
  • acq 196-199
  • spec 200-211
  • 9:57 seeing 0.9" on guider
  • 10:27 seeing 0.7" on guider
  • 11:00 some tiny cirrus at the end of the script

11:08 slewing to Prandoni USS172 imaging K
  • 11:18 seeing 0.5" on guider

11:49 executing Prandoni USS172 imaging J
  • error, the to check the pointing
  • 11:53 sending again, now is ok

12:14 slewing to Prandoni USS202 imaging K
  • 12:31 seeing 0.6" on guider

12:51 sending Prandoni USS202 imaging J

13:22 calibration for Saracco, flat and arc
  • flat 276-285 lamps off-on
  • arc 286- off/on
Topic revision: r6 - 19 Mar 2018, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback